# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v2

@@ 2.2 DPSGD with Double Clipping
[…]
@@ 3 Main Theoretical Results
In this section, we construct RDP bounds to analyze the privacy loss associated with releasing the final-round model of DPSGD-GC and DPSGD-DC, respectively. To complete our analysis, we also provide the convergence analysis and derive the utility bounds.
@@ 3.1 Privacy Analysis of DPSGD
First, we provide the assumption used throughout our analysis.
@@
($L$-smooth of stochastic loss function). The stochastic loss function $l_{\xi}(\cdot):\mathbb{R}^{d}\times\mathcal{D}\mapsto\mathbb{R}$ is smooth with constant $L>0$. That is, for any $\xi\in\mathcal{D}$, $l_{\xi}(\boldsymbol{\theta})$, the loss accrued on model $\boldsymbol{\theta}$, is continuously differentiable, and the gradient $\nabla l_{\xi}(\cdot)$ is Lipschitz continuous with constant $L$, i.e.,
$\left\|\nabla l_{\xi}(\boldsymbol{\theta})-\nabla l_{\xi}\left(\boldsymbol{\theta}^{\prime}\right)\right\|\leq L\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|,$ (9)
for all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{d}$.
Define the noisy update function in DPSGD as $\psi(\boldsymbol{\theta}_{t})=\boldsymbol{\theta}_{t}-\frac{\eta}{b}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}_{C}\left(\nabla l_{\xi}\left(\boldsymbol{\theta}_{t}\right)\right)+\boldsymbol{\varrho}_{t}$, where $\boldsymbol{\varrho}_{t}\sim\mathcal{N}(\boldsymbol{0},\beta\eta^{2}\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ denotes the partial zero-mean Gaussian perturbation with a given constant $\beta\in(0,1)$. Then, we divide the original Gaussian noise $\mathcal{N}(\boldsymbol{0},\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ into two parts: $\mathcal{N}(\boldsymbol{0},\beta\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ and $\mathcal{N}(\boldsymbol{0},(1-\beta)\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$. The first part together with the stochastic gradient descent constitute the noisy update function, and Lemma 3.2 presented shortly is used to measure its privacy. The other part aims for reducing the shift amount of the shifted Rényi divergence leveraging Lemma 2.6, and the privacy loss associated with it can be measured.
Next, we introduce a key lemma that provides an upper bound on the shifted Rényi divergence for noisy update with general smooth loss functions.
@@
(Noisy smooth-reduction). Let $\psi$ and $\psi^{\prime}$ be two noisy update functions of DPSGD based on adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. If the stochastic loss function is $L$-smooth (3.1), for any random variables $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, we have
$\leq$ $\mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\textup{DP}}^{2}}.$
If we further assume $b\leq\frac{n}{5}$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, then
$\leq$ $\mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{8\alpha C^{2}}{\beta n^{2}\sigma_{\textup{DP}}^{2}}.$
Proof sketch.
We summarize the proof steps as follows. First, we transform the shifted Rényi divergence to the standard Rényi divergence utilizing the smoothness of losses and equivalent definitions of $\infty$-Wasserstein distance. Next, the post-processing and partial convexity inequalities allow us to derive the privacy loss associated with SGD sampling. Finally, we apply the strong composition lemma (Lemma A.10) of RDP and obtain the privacy loss associated with two Gaussian distributions. The complete proof can be found in Section B.1. ∎
This result generalizes the contraction-reduction lemma (Lemma 2.7) and its variants (Altschuler & Talwar, 2022; Altschuler et al., 2024), which all rely on the convexity of the loss function to ensure that the update function is contractive. It characterizes the privacy dynamics of shifted Rényi divergence for noisy stochastic updates with non-convex and smooth loss functions. Based on this building block, we now present the privacy guarantees for DPSGD-GC and DPSGD-DC, respectively.
Theorem 3.3.
(Privacy guarantee for DPSGD-GC). Given any number of iterations $T$, dataset size $n$, batch size $b$, stepsize $\eta$, constant $\beta\in(0,1)$, $\alpha>1$, gradient clipping threshold $C$, and noise scale $\sigma_{\textup{DP}}$, if the stochastic loss function is $L$-smooth (3.1), then the DPSGD-GC algorithm satisfies $(\alpha,\varepsilon)$-RDP for
$\varepsilon=\frac{2\alpha C^{2}}{\beta nb\sigma_{\textup{DP}}^{2}}T.$ (12)
If we further assume $b\leq\frac{n}{5}$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, then
$\varepsilon=\frac{8\alpha C^{2}}{\beta n^{2}\sigma_{\textup{DP}}^{2}}T.$ (13)
Proof sketch.
We establish our result utilizing the induction hypothesis from $T$ to 00 and the flexibility of the shifted Rényi divergence. For the base case at $t=0$, we have $\mathcal{D}_{\alpha}^{(z_{0})}(\mathbb{P}_{\boldsymbol{\theta}_{0}}||\mathbb{P}_{\boldsymbol{\theta}_{0}^{\prime}})=0$ since the initializations satisfy $\boldsymbol{\theta}_{0}=\boldsymbol{\theta}^{\prime}_{0}$. For the inductive step, we apply the noisy smooth-reduction lemma (Lemma 3.2) and subsequently reduce the shift amount using shift-reduction lemma (Lemma 2.6) with auxiliary variables to derive the recurrence relationship. Finally, by tracking the privacy loss across all iterations, we derive an upper bound on the RDP loss. The complete proof can be found in Section B.2. ∎
Theorem 3.4.
(Privacy guarantee for DPSGD-DC). Given any number of iterations $T$, dataset size $n$, batch size $b$, stepsize $\eta$, constant $\beta\in(0,1)$, $\alpha>1$, gradient clipping threshold $C$, diameter of parameter domain $D$, and noise scale $\sigma_{\textup{DP}}$, if the stochastic loss function is $L$-smooth (3.1), then the DPSGD-GC algorithm satisfies $(\alpha,\varepsilon)$-RDP for
$\varepsilon=\frac{2\alpha C^{2}}{\beta nb\sigma_{\textup{DP}}^{2}}+\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\textup{DP}}^{2}(1-\beta)}.$ (14)
If we further assume $b\leq\frac{n}{5}$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, then
$\varepsilon=\frac{8\alpha C^{2}}{\beta n^{2}\sigma_{\textup{DP}}^{2}}+\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\textup{DP}}^{2}(1-\beta)}.$ (15)
Proof sketch.
This proof follows a similar approach to Theorem 3.3 but terminates the induction early at time step $\tau$. By judiciously setting the shift amount $z_{\tau}$ as $D$ at $t=\tau$, we obtain $\mathcal{D}_{\alpha}^{(z_{\tau})}(\mathbb{P}_{\boldsymbol{\theta}_{\tau}}||\mathbb{P}_{\boldsymbol{\theta}_{\tau}^{\prime}})=0$ due to the bounded domain assumption. Finally, by appropriately configuring the values of the auxiliary shift variables, we derive the converged privacy loss under the bounded domain restriction, as detailed in Section B.3. ∎
Remark 3.5.
Comparing Theorem 3.3 and Theorem 3.4, we observe that the privacy loss can still converge to a constant for non-convex smooth losses when the domain is bounded by $D$. However, unlike the convex case (Altschuler & Talwar, 2022; Altschuler et al., 2024), where the upper bound scales linearly with $D$, the non-convex setting results in a bound that scales quadratically with $D$. This suggests that privacy bounds in non-convex scenarios are inherently looser than those in convex cases, consistent with recent findings (Kong & Ribero, 2024; Chien & Li, 2024).
3.2 Utility Analysis of DPSGD
Next, we provide the convergence bounds (utility) for DPSGD-GC and DPSGD-DC under $(\alpha,\varepsilon)$-RDP constraint, respectively. All bounds are expressed as expectations over the randomness of SGD and Gaussian noise. Our utility analysis builds upon the results of Bassily et al. (2014), Zhang et al. (2020), and Koloskova et al. (2023).
Assumption 3.6.
(Bounded SGD variance). The stochastic gradient $\nabla l_{\xi}$ has bounded variance, that is, for all $\boldsymbol{\theta}\in\mathbb{R}^{d}$, we have
Note that 3.1 implies that the loss function $l(\cdot)=\mathbb{E}_{\xi}[l_{\xi}(\cdot)]$ is also smooth with constant $L>0$. Based on 3.1 and 3.6, the following lemma provides an upper bound on the minimum expected norm of the gradient for DPSGD-GC with smooth loss functions.
Lemma 3.7.
@@ 3.2 Utility Analysis of DPSGD
(Convergence result of DPSGD-GC (Koloskova et al., 2023)). Assume that the loss function $l(\cdot)$ has smoothness parameter $L$ and SGD variance at most $\sigma_{\textup{SGD}}^{2}$ (3.6). When running DP-SGD-GC for $T$ steps with step-size $\eta\leq\frac{1}{9L}$, the minimum expected norm of the gradient $\min_{t\in[0,T]}\mathbb{E}[\|\nabla l(\boldsymbol{\theta}_{t})\|]$ is upper bounded by
@@
($\mu$-strongly convex). The loss function $l(\cdot)$ is strongly convex with constant $\mu>0$ if and only if the following inequality holds
$\left[\nabla l(\boldsymbol{\theta})-\nabla l\left(\boldsymbol{\theta}^{\prime}\right)\right]^{\top}\left(\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right)\geq\mu\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|^{2},$ (18)
for all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{d}$.
For DPSGD-DC, we focus on smooth and strongly convex losses and obtain the following result.
@@
(Convergence result of DPSGD-DC). Assume that the loss function $l(\cdot)$ has smoothness parameter $L$, strongly convex parameter $\mu$ (3.8), SGD variance at most $\sigma_{\textup{SGD}}^{2}$ (3.6), and $\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}l(\boldsymbol{\theta})\in\operatorname{int}\mathcal{K}$. When running DPSGD-DC for $T$ steps with step-size $\eta\leq\frac{9}{20L}$, the convergence in terms of $\min_{t\in[0,T]}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]$ is upper bounded by
@@
$\leq\mathcal{O}\biggl{(}\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{D}{\sqrt{\eta T}}+\min\bigl{(}\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\textup{SGD}},\sqrt{\frac{\sigma_{\textup{SGD}}^{3}}{\mu C}}\bigr{)}$
$+\frac{\sqrt{\eta}\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{d\eta\sigma_{\textup{DP}}^{2}\sqrt{L}}{C}+\sqrt{d\eta}\sigma_{\textup{DP}}\biggr{)},$
@@
We present an intuitive proof sketch below, with the complete proof provided in Section C.1. The primary challenges stem from the gradient clipping operation, the SGD procedure, and the parameter projection step. To address these, we divide the proof into separate cases. For instance, in the case where the clipping threshold $C\leq 10\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{2}}$ and $\|\nabla l(\boldsymbol{\theta}_{t})\|\geq 35\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{3}{4}}$, first, we leverage the non-expansiveness of the projection operator to analyze the impact of the parameter projection. Next, we introduce an auxiliary clipping factor $\gamma_{\xi}=\min(1,\frac{C}{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})\|})$ and apply the Markov inequality to quantify the effects of clipped SGD. Through careful step-size design and manipulations, we obtain a recurrence relation that characterizes the evolution over two successive time steps. Finally, by averaging over $t$, we derive an upper bound for this case. The proof for other cases follows a similar structure but incorporates different auxiliary variables tailored to the true gradient. By summing up all cases, we derive an upper bound on the error for DPSGD-DC. ∎
@@
Note that our results use a different convergence metric, $\min_{t\in[0,T]}\mathbb{E}[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}]$, instead of the more commonly used $\min_{t\in[0,T]}\mathbb{E}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$, as it better facilitates our derivation and gaining insights into the analytical results.
Using our RDP guarantees in Theorem 3.3 and Theorem 3.4, we immediately obtain the following results.
Corollary 3.11.
(Privacy-utility trade-off for DPSGD-GC). Assuming that the conditions in Lemma 3.7 are satisfied, for $\boldsymbol{\theta}_{T}$ output by DPSGD-GC with $L$-smooth losses, we have the following results.
  • For $\sigma_{\textup{DP}}^{2}=\mathcal{O}\bigl{(}\frac{\alpha C^{2}T}{\varepsilon nb}\bigr{)}$, we have the following inequality
    $+\sqrt{\eta L}\frac{\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{d\alpha CL\eta T}{\varepsilon nb}+\frac{\sqrt{d\alpha L\eta T}C}{\sqrt{\varepsilon nb}}\biggr{)}.$
  • If we further assume $b\leq\frac{n}{5}$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, and let $\sigma_{\textup{DP}}^{2}=\mathcal{O}\bigl{(}\frac{\alpha C^{2}T}{\varepsilon n^{2}}\bigr{)}$, then we have the following inequality
    $+\sqrt{\eta L}\frac{\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{d\alpha CL\eta T}{\varepsilon n^{2}}+\frac{\sqrt{d\alpha L\eta T}C}{\sqrt{\varepsilon}n}\biggr{)}.$
Corollary 3.12.

### v3

@@ 2.2 DPSGD with Double Clipping (DPSGD-DC)
  3) update and project the parameters: $\boldsymbol{\theta}_{t+1}\leftarrow\Pi_{\mathcal{K}}(\boldsymbol{\theta}_{t}-\eta(\nabla\mathcal{L}_{B_{t}}\left(\boldsymbol{\theta}_{t};\mathcal{D}\right)+\boldsymbol{\zeta}_{t}))$,where $\Pi_{\mathcal{K}}(\boldsymbol{\theta})\!=\!\arg\min_{\boldsymbol{x}\in\mathcal{K}}\|\boldsymbol{\theta}-\boldsymbol{x}\|$ and $\boldsymbol{\zeta}_{t}\sim\!\mathcal{N}(\boldsymbol{0},\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$;
@@
In this paper, we also consider the DPSGD with both gradient clipping and parameter projection (Algorithm 1), termed as DPSGD-DC. This method begins with applying the vanilla SGD procedure using gradient clipping and Gaussian perturbation for model updates. Then, the updated parameters are projected into a bounded domain $\mathcal{K}=\{\boldsymbol{\theta}\in\mathbb{R}^{d}:\|\boldsymbol{\theta}\|\leq D\}$ as follows [^2] A detailed discussion of the bounded domain assumption is provided in Appendix E.1.
with
where $\eta$ denotes the learning rate; $\boldsymbol{\zeta}_{t}\sim\mathcal{N}(\boldsymbol{0},\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ denotes a multi-variate Gaussian random variable at iteration $t$; the loss function accrued on a model using data sample $\xi\in\mathcal{D}$ is defined as $l_{\xi}(\cdot):\mathcal{D}\times\mathbb{R}^{d}\mapsto\mathbb{R};$ and $\nabla\mathcal{L}_{B_{t}}\left(\boldsymbol{\theta}_{t};\mathcal{D}\right)$ is the clipped gradient of the loss function averaging over a mini-batch $\mathcal{B}_{t}$, i.e., $\nabla\mathcal{L}_{B_{t}}\left(\boldsymbol{\theta}_{t};\mathcal{D}\right)=\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}_{C}(\nabla l_{\xi}(\boldsymbol{\theta}_{t}))$, with $b=\left|\mathcal{B}_{t}\right|$ denoting the size of the mini-batch and $\operatorname{clip}_{C}(\boldsymbol{x})=\boldsymbol{x}\cdot\min(1,\frac{C}{\|\boldsymbol{x}\|})$. Note that if $\mathcal{K}=\mathbb{R}^{d}$, it reduces to DPSGD with only gradient clipping, i.e., DPSGD-GC.
@@ 3 Main Theoretical Results
In this section, we construct RDP bounds to analyze privacy loss associated with releasing only the final-round model of DPSGD-GC and DPSGD-DC, respectively. To gain insights into how such privacy loss affects their respective training performance, we further provide utility analysis and derive the corresponding privacy-utility trade-offs.
Assumption 3.1.
($L$-smoothness of the loss function). The loss function $l_{\xi}(\cdot):\mathcal{D}\times\mathbb{R}^{d}\mapsto\mathbb{R}$ is smooth with constant $L>0$, if for any $\xi\in\mathcal{D}$ and all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{d}$, $l_{\xi}(\boldsymbol{\theta})$ is continuously differentiable, and its gradient $\nabla l_{\xi}(\cdot)$ in terms of $\boldsymbol{\theta}$ is $L$-Lipschitz, i.e.,
$\left\|\nabla l_{\xi}(\boldsymbol{\theta})-\nabla l_{\xi}(\boldsymbol{\theta}^{\prime}\right)\|\leq L\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|.$ (8)
@@ 3.1 Privacy Analysis of DPSGD
In this subsection, first, we divide the original additive Gaussian noise $\boldsymbol{\zeta}_{t}\sim\mathcal{N}(\boldsymbol{0},\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ (c.f. (6)) into two parts: $\boldsymbol{\varrho}_{t}\sim\mathcal{N}(\boldsymbol{0},\beta\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ and $\boldsymbol{\varsigma}_{t}\sim\mathcal{N}(\boldsymbol{0},(1-\beta)\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$. The first part, together with the clipped SGD update, constitutes the noisy update function, defined as
$\psi(\boldsymbol{\theta}_{t})\stackrel{{\scriptstyle\Delta}}{{=}}\boldsymbol{\theta}_{t}-\eta\left(\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}_{C}\left(\nabla l_{\xi}\left(\boldsymbol{\theta}_{t}\right)\right)+\boldsymbol{\varrho}_{t}\right),$ (9)
and the privacy loss of which is characterized by Lemma 3.2 presented shortly; the other part, $\eta\boldsymbol{\varsigma}_{t}$, aims for reducing the shift amount of the shifted Rényi divergence leveraging Lemma 2.6.
Next, we provide the following key lemma that provides the upper bound on the shifted Rényi divergence between the distributions of the noisy update function applied on two adjacent datasets with a smooth loss function.
Lemma 3.2.
(Noisy smooth-reduction). Let $\psi(\cdot)$ and $\psi^{\prime}(\cdot)$ be noisy update functions (c.f. (9)) of DPSGD based on adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, respectively, and $n$ be the size of $\mathcal{D}$. If the loss function is $L$-smooth (Assumption 3.1), for any random variables $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, we have
$\mathcal{D}_{\alpha}^{((1+\eta L)z)}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})\leq\ \mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\textup{DP}}^{2}}.$ (10)
Proof sketch.
We summarize the proof steps as follows. First, we transform the shifted Rényi divergence to the standard Rényi divergence utilizing the smoothness of losses and equivalent definitions of $\infty$-Wasserstein distance. Next, the post-processing (Lemma A.9) and partial convexity inequality (Lemma A.10) allow us to derive the privacy loss associated with SGD sampling. Finally, we apply the strong composition lemma (Lemma A.11) of RDP, and obtain the privacy loss associated with these two Gaussian distributions. The complete proof can be found in Appendix B.1. ∎
If we further assume the mini-batch size $b\leq\frac{n}{5}$, the RDP parameter $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and the per-dimension Gaussian noise scale $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, a strengthened result can be obtained. Similar findings hold for all of our following theoretical results, and the complete version of Lemma 3.2 can be found in Appendix B.1.
Note that this result generalizes the contraction-reduction lemma (see Lemma A.1) (feldman2018privacy) and its variants (altschuler2022privacy; altschuler2024privacy) in existing literature, which all rely on the convexity of the loss function to ensure that the update function is contractive. [^3] A function is said to be contractive if it is $1$-Lipschitz. It characterizes the privacy dynamics of shifted Rényi divergence for noisy stochastic updates with general smooth loss functions. Based on this building block, we are ready to present the privacy guarantees for DPSGD-GC and DPSGD-DC, respectively.
Theorem 3.3.
(Privacy guarantee for DPSGD-GC). Given the number of total iterations $T$, dataset size $n$, batch size $b$, stepsize $\eta$, $\alpha>1$, gradient-clipping threshold $C$, and noise scale $\sigma_{\textup{DP}}$, if the loss function is $L$-smooth (Assumption 3.1), then the DPSGD-GC algorithm satisfies $(\alpha,\varepsilon)$-RDP for
$\varepsilon=\mathcal{O}\left(\frac{\alpha C^{2}}{nb\sigma_{\textup{DP}}^{2}}T\right).$ (11)
Proof sketch.
We establish our result utilizing the recursive hypothesis from $T$ to 0 and the flexibility of the shifted Rényi divergence. For the base case at $t=0$, we have $\mathcal{D}_{\alpha}^{(z_{0})}(\mathbb{P}_{\boldsymbol{\theta}_{0}}||\mathbb{P}_{\boldsymbol{\theta}_{0}^{\prime}})=0$, since the initialization satisfy $\boldsymbol{\theta}_{0}=\boldsymbol{\theta}^{\prime}_{0}$. For the recursive step, we apply Lemma 3.2 and subsequently reduce Lemma 2.6 with auxiliary variables to derive the recurrence relationship. Finally, by tracking the privacy loss across all iterations, we derive an upper bound on the RDP loss. The complete proof can be found in Appendix B.2. ∎
Theorem 3.4.
(Privacy guarantee for DPSGD-DC). Given the number of total iterations $T$, dataset size $n$, batch size $b$, stepsize $\eta$, $\alpha>1$, gradient-clipping threshold $C$, diameter of the model-parameter domain $D$, and noise scale $\sigma_{\textup{DP}}$, if the loss function is $L$-smooth (Assumption 3.1), then the DPSGD-DC algorithm satisfies $(\alpha,\varepsilon)$-RDP for
$\varepsilon=\mathcal{O}\left(\frac{\alpha C^{2}}{nb\sigma_{\textup{DP}}^{2}}\min\left\{T,\frac{(1+\eta L)^{2}nbD^{2}}{\eta^{2}C^{2}}\right\}\right).$ (12)
Proof sketch.
This proof follows a similar approach to Theorem 3.3 but terminates the recursion early at iteration $\tau$. By judiciously setting the shift amount $z_{\tau}=D$ at $t=\tau$, we obtain $\mathcal{D}_{\alpha}^{(z_{\tau})}(\mathbb{P}_{\boldsymbol{\theta}_{\tau}}||\mathbb{P}_{\boldsymbol{\theta}_{\tau}^{\prime}})=0$ due to the bounded domain assumption. Finally, by appropriately setting the values of the auxiliary shift variables, we derive the converged privacy loss, as detailed in Appendix B.3. ∎
Remark 3.5.
Comparing Theorem 3.4 to 3.3, we observe that the privacy loss for DGSGD-DC can still converge to a constant for non-convex smooth loss function when the model-parameter domain is bounded by $D$. However, unlike the convex cases (altschuler2022privacy; altschuler2024privacy), where the upper bound scales linearly with $D$, the non-convex setting yields a bound that scales quadratically with $D$. This suggests that upper bounds on RDP for bounded-domain scenarios are inherently looser without the assumption on convexity, which is consistent with intuition and recent findings (kong2024privacy; chien2024convergent).
3.2 Utility Analysis of DPSGD
In this subsection, we provide upper bounds on utility functions that reflect convergence behavior of DPSGD-GC and DPSGD-DC under $(\alpha,\varepsilon)$-RDP constraint, respectively. All bounds are expressed as expectations over the randomness of SGD sampling and Gaussian noise.
@@ 3.2 Utility Analysis of DPSGD
(Bounded SGD variance). The gradient $\nabla l_{\xi}(\cdot)$ has bounded variance. That is, for all $\boldsymbol{\theta}\in\mathbb{R}^{d}$, we have
Note that Assumption 3.1 implies that the population risk function $l(\cdot)=\mathbb{E}_{\xi}[l_{\xi}(\cdot)]$ is also smooth with constant $L>0$. Based on Assumption 3.1 and 3.6, the following lemma provides an upper bound on the minimum expected norm of the gradient for DPSGD-GC with smooth population risk functions.
@@
(Utility bound for DPSGD-GC (koloskova2023revisiting)). Assume that the population risk function $l(\cdot)$ has smoothness parameter $L$ and SGD variance bounded from above by $\sigma_{\textup{SGD}}^{2}$ (Assumption 3.6). The DPSGD-GC running for $T$ iterations with step-size $\eta\leq\frac{1}{9L}$ has the minimum expected norm of the population risk’s gradient upper bounded by
@@
Proposition 3.8.
(Privacy-utility trade-off for DPSGD-GC). Assuming that the conditions in Lemma 3.7 are satisfied, for DPSGD-GC with $L$-smooth population risk and $\sigma_{\textup{DP}}^{2}=\mathcal{O}\bigl(\frac{\alpha C^{2}T}{\varepsilon nb}\bigr)$, we have the following results:
$\leq\mathcal{O}\biggl(\frac{1}{\eta CT}+\frac{1}{\sqrt{\eta T}}+\min\left\{\sigma_{\textup{SGD}},\frac{\sigma_{\textup{SGD}}^{2}}{C}\right\}$
$+\sqrt{\eta L}\frac{\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{\alpha dL\eta CT}{\varepsilon nb}+\frac{\sqrt{\alpha dL\eta T}C}{\sqrt{\varepsilon nb}}\biggr).$
Following this proposition, we can derive the achievable privacy-utility trade-off for DPSGD-GC as $\mathcal{O}(\max\{\frac{dL\log(1/\delta)}{\epsilon^{2}n^{2}},\sigma_{\text{SGD}}^{4/3}[\frac{dL\log(1/\delta)}{\epsilon^{2}n^{2}}]^{1/3}\})$, and we defer the details to Appendix C.2.
@@
($\mu$-strongly convex). The population risk function $l(\cdot)$ is strongly convex with constant $\mu>0$ if and only if the following inequality holds for all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{d}$, i.e.,
$\left[\nabla l(\boldsymbol{\theta})-\nabla l\left(\boldsymbol{\theta}^{\prime}\right)\right]^{T}\left(\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right)\geq\mu\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|^{2}.$ (16)
For DPSGD-DC, in the special case of smooth and strongly convex population risk, we obtain the following upper bound that describes the minimum expected square root of the optimality gap.
@@
(Utility bound for DPSGD-DC). Assume that the population risk function $l(\cdot)$ has smoothness parameter $L$, strongly convex parameter $\mu$ (Assumption 3.9), SGD variance bounded from above by $\sigma_{\textup{SGD}}^{2}$ (Assumption 3.6), and $\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}l(\boldsymbol{\theta})\in\operatorname{int}\mathcal{K}$. The DPSGD-DC running for $T$ iterations with step-size $\eta\leq\frac{9}{20L}$ has a minimum expected square root of the optimality gap upper bounded by
@@
$\leq\mathcal{O}\biggl(\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{D}{\sqrt{\eta T}}+\min\left\{\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\textup{SGD}},\sqrt{\frac{\sigma_{\textup{SGD}}^{3}}{\mu C}}\right\}$
$+\frac{\sqrt{\eta}\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{d\eta\sigma_{\textup{DP}}^{2}\sqrt{L}}{C}+\sqrt{d\eta}\sigma_{\textup{DP}}\biggr).$
@@
We present an intuitive sketch below, with the complete proof provided in Appendix C.1. The primary challenges stem from the gradient clipping operation, the SGD procedure, and the parameter projection step. To address these, we divide the proof into several cases. For instance, in the case where the clipping threshold $C\leq 10\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{2}}$ and $\|\nabla l(\boldsymbol{\theta}_{t})\|\geq 35\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{3}{4}}$, first, we leverage the non-expansiveness of the projection operator to analyze the impact of the parameter projection. Next, we introduce an auxiliary clipping factor $\gamma_{\xi}=\min(1,\frac{C}{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})\|})$ and apply the Markov inequality to quantify the effects of clipped SGD. By careful step-size design and manipulations, we obtain a recurrence relation that characterizes the evolution over two successive time steps. Finally, by averaging over $t$, we derive an upper bound for this case. The proof for other cases follows a similar procedure but incorporates different auxiliary variables tailored to the true gradient. By summing up all cases, we derive an upper bound on the error for DPSGD-DC. ∎
@@
Note that our results adopt a different utility metric other than the more commonly used $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$, as this metric better facilitates our analysis and gaining insights into the results.
Using the RDP guarantees in Theorem 3.4, we immediately obtain the following results.
Proposition 3.12.